OpenStudy (anonymous):

a student runs 100 meters in 11 seconds. What is the speed of the student? speed=distance/time; 1 mile= 1609 meters

OpenStudy (natalieee):

you gave yourself the answer . fill in the equation with the numbers you have .

OpenStudy (texaschic101):

s = d/t s = 100/11 = 9.09 meters per second